5. The stock market A stock market isa market for trading a companys stocks and derivatives. In a dealer market, some dealers hold a oertain inventory of specific securities and create a liquid market by purchasing and selling their inventories. These dealers make a market and are thus called market makers. Agents in the market bring investors to the dealers through a network of terminals and electronic systems. Where do dealer profits come from in a dealer market? O Dividend payments O The bid-ask spread O Interest charges Brent, a trader, wants to buy 1,ODD shares of XYZ stock, while a second trader, Amara, is willing to sell 1,500 shares of the same stock. Unfortunately, Brent and Amara dant know one anather, and must complete their transactions using the stock exchanges market-making dealer. XYZs market-maker is willing to sell her shares for $28.55 per share and purchase additional shares for $27.25 per share. Select the most appropriate values in the following table Term Value Bid price Ask price Bid-ask spread If the market-maker is willing to purchase the entire black of 1,500 shares from Amara and, from that block, resell 1,000 shares to Brent, then the market-makers net profit from Brents transaction-excluding any inventory effects- will be Consider this case: Last year Haywoad Inc. entered into a firm commitment agreement with Barday Partners, an investment bank. At the time of issue, Barday Partners has agreed to purchase all offered shares from Haywood and then try to sell all shares in the primary market. In the event that an issuer elects to use an underwritben arrangement, who bears the entire risk that the stock issue might be undersubscribed? In other words, who is at risk if the investment bank cannot sell all shares to investors at the time of issue? The investment bank O The issuing firm
Answer #1

1: Public

The stock market is open to the public for trading

2: The bid ask spread

Dealers buy at a lower price and sell at a higher price. This is called the bid ask spread and is the dealers’ profit.

3: Bid price = $27.25 (Purchase price)

Ask price = $28.55 (Selling price)

Bid ask spread = 28.55- 27.25 = $1.3

4: Profit = Bid ask spread* Quantity sold

= 1.3* 1000

=$1300

5: The investment bank

Since this is an underwriting arrangement the investment bank has to buy all the unsold shares.